OpenStudy (anonymous):

how far is the sears tower [at point (-6,-3)] from Navy Pier [at point (10,9)], given that each unit is approximately 1/8 of a mile

OpenStudy (anonymous):

i have already calculated the distance in units is 20, i am confused about how to transfer it to miles

OpenStudy (jdoe0001):

well, 20 units times 1/8 should give you how many miles for example if you only had 8 units 8 * 1/8 = 8/8 = 1mile

OpenStudy (anonymous):

so 2.5 miles

OpenStudy (jdoe0001):

1/8 + 1/8 + 1/8 + 1/8 + 1/8 +1/8 +1/8+1/8 = 1mile

OpenStudy (jdoe0001):

2 miles and a half, yes